OpenStudy (anonymous):

What is the solution to this equation? 10x - 15x + 5 = -45 + 20 A. x = 6 B. x = 4 C. x = -6 D. x = -4

OpenStudy (anonymous):

-5x=-30 x=6

OpenStudy (anonymous):

A

OpenStudy (anonymous):

10x - 15x +5 = -45 + 20 (bring like terms together) 10x - 15x = -45 + 20 - 5 -5x = -30 x = -30/-5 x=6